# DriveDreamer4D: World Models Are Effective Data Machines for 4D Driving Scene Representation

### v2

@@ 4.2 Comparison with Different 4DGS Baselines
Qualitative Results. In addition to quantitative comparisons, we provide a qualitative analysis of novel trajectory view renderings, as shown in Fig. 3, focusing on speed change scenarios. Our method significantly enhances the positional accuracy of foreground vehicles and background elements under acceleration viewpoints. Specifically, baseline algorithm results (PVG [8], $\text{S}^{3}$aussian [26], Deformable-GS [69]) are displayed in rows 1, 3, and 5. It is clear that the baseline methods struggle with perspective synthesis under acceleration, leading to inaccurate positional shifts of surrounding vehicles during ego-car acceleration. In contrast, with the integration of DriveDreamer4D, the 4DGS algorithms achieve both enhanced spatial consistency and noticeably improved rendering quality. Especially noticeable in the last frame (rightmost column of Fig. 3), the baseline algorithms lead to blurring or disappearance of foreground vehicles. In contrast, our method significantly improves image rendering quality, as highlighted by the orange boxes. In Fig. 4, we present the novel trajectory view synthesis during lane change. Images rendered by the baseline algorithms exhibit issues where foreground vehicles incorrectly change lanes in sync with the camera’s motion, and some vehicles are incompletely rendered. Additionally, the background is filled with speckles and ghosting. Especially shown in the last frame (rightmost column of Fig. 4), baseline algorithms often produce blurred, ghosted foreground vehicles and background speckles in the sky, alongside blurred lane markings. Our method, however, significantly improves rendering quality, as highlighted by the orange boxes. Vehicle contours are sharper, and background artifacts such as speckles and ghosting are substantially reduced.
4.3 Ablation Studies
We conduct an ablation study based on PVG [8] to determine the optimal mixing ratio $\beta$ for combining real and synthetic data. As shown in Tab. 5, our results indicate that increasing the proportion of generated data ($\beta=0.2\sim 0.5$) significantly improves both NTA-IoU and FID metrics compared to using only real data ($\beta=0$). To balance the FID and NTA-IoU metrics, we ultimately selected $\beta=0.4$. Additionally, an ablation study is performed to identify the appropriate training loss weight $\lambda_{\text{novel}}$ for the novel trajectory view. Results presented in Tab. 5 demonstrate that incorporating the loss for the novel trajectory view while optimizing 4DGS algorithms effectively enhances the NTA-IoU and FID metrics. Taking into account the balance between the FID and NTA-IoU scores, the $\lambda_{\text{novel}}$ is set to 1. Finally, as shown in Table 6, the experiment confirms that the depth loss should not be included when optimizing novel trajectory views, since LiDAR depth maps are incomplete due to occlusions.
$\beta$ NTA-IoU $\uparrow$ FID $\downarrow$
0 0.351 105.29
0.2 0.413 82.37
0.3 0.417 80.68
0.4 0.420 79.54
0.5 0.425 83.19
Table 4: Ablation study on the mixing ratio $\beta$.
@@ 4.3 Ablation Studies
Table 5: Ablation study on the training loss weight $\lambda_{\text{novel}}$.
depth loss NTA-IoU $\uparrow$ FID $\downarrow$
$\checkmark$ 0.401 82.63
$\times$ 0.420 79.54
Table 6: Ablation study on the depth loss of the generated data.
@@ 5 Discussion and Conclusion
In this paper, we presented DriveDreamer4D, a novel framework designed to advance 4D driving scene representations by harnessing priors from world models. Addressing key limitations of current sensor simulation methods—namely, their dependence on forward-driving training data distributions and inability to model complex maneuvers—DriveDreamer4D leverages a world model to generate novel trajectory videos that complement real-world driving data. By explicitly employing structured conditions, our framework maintains spatial-temporal consistency across foreground and background elements, ensuring that generated data adheres closely to the dynamics of real-world traffic scenarios. Our experiments demonstrate that DriveDreamer4D achieves superior quality in generating diverse simulation viewpoints, with significant improvements in both the rendering fidelity and spatiotemporal coherence of scene components. Notably, these results highlight DriveDreamer4D’s potential as a foundation for closed-loop simulations that require high-fidelity reconstructions of dynamic driving scenes.
Acknowledgments
The authors would like to thank Chen Liu, Yuyin Chen, Kun Zhan, Peng Jia, Yifei Zhan, and Fu Liu at Li Auto Inc. for the fruitful discussions, related implementation, and the contributed computational resources for this work.

### v3

@@ 4.1 Experiment Setup
4.2 Comparison with Different 4DGS Baselines
Quantitative Results. As demonstrated in Tab. 1, integrating DriveDreamer4D with different 4DGS algorithms consistently yields superior NTA-IoU and NTL-IoU scores across diverse, complex maneuvers (e.g., lane changes, acceleration, and deceleration), significantly outperforming the baseline methods. Specifically, with DriveDreamer4D, the average NTA-IoU scores across three baselines (PVG [9], $\text{S}^{3}$aussian [28], Deformable-GS [74]) are relatively enhanced by 22.6%, 43.5%, and 15.6%, underscoring DriveDreamer4D’s capability to improve the spatiotemporal coherence of foreground agents. Moreover DriveDreamer4D facilitates a relative improvement in average NTL-IoU for these baselines by 1.8%, 3.7%, and 1.6%, thereby markedly enhancing the spatiotemporal coherence of background lanes in 4D rendering of driving scenarios.
In addition to verifying the spatiotemporal consistency of rendered novel trajectory views, we leverage the FID metric to assess rendering quality under novel trajectories. Given that acceleration and deceleration scenarios yield rendered views with distributional similarities to ground truth, limiting FID’s discriminative capability across algorithms, our FID comparisons focus specifically on lane change scenarios. Experiment results, as presented in Tab. 2, indicate that our method substantially outperforms the baseline methods (PVG [9], $\text{S}^{3}$aussian [28], Deformable-GS [74]), with FID relative improvements of 32.1%, 46.4%, and 16.3%. These results highlight DriveDreamer4D’s capability to enhance generation quality for novel trajectory viewpoints.
Finally, we conduct a user study to evaluate the rendering quality of different methods on novel trajectories, with a specific focus on foreground agents. For each method, we generate three novel trajectory views—lane change, acceleration, and deceleration—across eight scenes from the Waymo dataset [57]. Participants are then asked to select the renderings they found most visually favorable in each comparison. The DriveDreamer4D win rates from this study, shown in Tab. 3, reveal a significant user preference for our method’s renderings.
@@ 4.2 Comparison with Different 4DGS Baselines
Table 4: Ablation study on the training loss weight $\lambda_{\text{novel}}$ for novel trajectory data.
$\lambda_{\text{reg}}$ NTA-IoU $\uparrow$ FID $\downarrow$
0 0.420 79.54
1e-2 0.411 119.39
1e-3 0.428 71.52
1e-4 0.422 75.31
Table 5: Ablation study on the regularization loss weight $\lambda_{\text{reg}}$.
Depth Loss Temporal-aligned Cousin Pair Regularization Loss NTA-IoU $\uparrow$ FID $\downarrow$
$\checkmark$ $\times$ $\times$ 0.401 82.63
$\times$ $\times$ $\times$ 0.420 79.54
$\times$ $\checkmark$ $\times$ 0.423 76.20
$\times$ $\checkmark$ $\checkmark$ 0.428 71.52
Table 6: Ablation study on the depth loss and CDTS.
Qualitative Results. In addition to quantitative comparisons, we provide a qualitative analysis of novel trajectory view renderings. As shown in Fig. 3, we present the novel trajectory view synthesis during lane change. Images rendered by the baseline algorithms exhibit issues where foreground vehicles incorrectly change lanes in sync with the camera’s motion, and some vehicles are incompletely rendered. Additionally, the background is filled with speckles and ghosting. Especially shown in the rightmost column of Fig. 3, baseline algorithms often produce blurred, ghosted foreground vehicles and background speckles in the sky, alongside blurred lane markings. Our method, however, significantly improves rendering quality, as highlighted by the orange boxes. Vehicle contours are sharper, and background artifacts such as speckles and ghosting are substantially reduced.
4.3 Ablation Studies
We conduct an ablation study based on PVG [9], analyzing the effects of hyperparameters settings for $\lambda_{\text{novel}}$ and $\lambda_{\text{reg}}$, as well as the impact of depth loss, temporal-aligned cousin pairs, and regularization loss. Following the experiments shown in Tab. 5 and 5, we set $\lambda_{\text{novel}}$ to 1 and $\lambda_{\text{reg}}$ to $1\times 10^{-3}$, which yield the best performance. As shown in Tab. 6, the experiment confirms that the depth loss should be excluded when optimizing novel trajectory views, since LiDAR depth maps are incomplete due to occlusions. Furthermore, employing temporal-aligned cousin pairs achieves an FID of 76.20 and an NTA-IoU of 0.423, with the perception regularization loss further improving the FID to 71.52 and NTA-IoU to 0.428. These results highlight the effectiveness of CDTS, in achieving $\sim$10improvement in FID, along with a $\sim$2increase in NTA-IoU.
@@ 5 Discussion and Conclusion
In this paper, we presented DriveDreamer4D, a novel framework designed to advance 4D driving scene representations by harnessing priors from world models. Addressing key limitations of current sensor simulation methods—namely, their dependence on forward-driving training data distributions and inability to model complex maneuvers—DriveDreamer4D leverages a world model to generate novel trajectory videos that complement real-world driving data. By explicitly employing structured conditions, our framework maintains spatial-temporal consistency across traffic elements, ensuring that generated data adheres closely to the dynamics of real-world traffic scenarios. Our experiments demonstrate that DriveDreamer4D achieves superior quality in generating diverse simulation viewpoints, with significant improvements in both the rendering fidelity and spatiotemporal coherence of scene components. Notably, these results highlight DriveDreamer4D’s potential as a foundation for closed-loop simulations that require high-fidelity reconstructions of dynamic driving scenes.